# Optimal Computation in Anonymous Dynamic Networks

### v7

@@ 6 Negative Results
In this section we collect several negative results and counterexamples, which not only provide lower bounds asymptotically matching our algorithms’ running times, but also justify all of the assumptions made in Sections 4 and 5 about the a-priori knowledge of processes. Although some of these facts were previously known, here we offer simple and self-contained proofs based on history trees.
@@ 6.1 Unsolvable Problems
Showing that certain problems are not solvable in certain network models allows us to argue that our algorithms are universal, i.e., they can be applied to all solvable problems. Remarkably, all of our algorithms exhibit the highest level of generality possible.
Networks with leaders. We will prove that the multiset-based functions introduced in Section 2 are the only functions that can be computed deterministically in anonymous networks with leaders. This result, together with Proposition 2.1, shows that the algorithms in Sections 4.4 and 5 can be generalized to all functions that can be computed in networks with leaders, i.e., the multiset-based functions.
@@
No function other than the multiset-based functions can be computed (with or without termination), even when restricted to connected static simple networks with a known and arbitrary number of leaders.
@@
Let us consider the (static) network whose topology at round $t$ is the complete graph $G_{t}=K_{n}$, i.e., each process receives messages from all other processes at every round. We can prove by induction that all nodes of the history tree other than the root have exactly one child. This is because any two processes with the same input always receive equal multisets of messages, and are therefore always indistinguishable. Thus, the history tree is completely determined by the multiset $\mu_{\lambda}$ of all processes’ inputs; moreover, a process’ view at any given round only depends on the process’ own input and on $\mu_{\lambda}$. By the fundamental theorem of history trees Theorem 3.1, this is enough to conclude that if a process’ output stabilizes, that output must be a function of the process’ own input and of $\mu_{\lambda}$, which is the defining condition of a multiset-based function. ∎
The following result justifies the assumption made in Sections 4.4 and 5 that processes have a-priori knowledge of the number of leaders $\ell$ in the system.
@@
No algorithm can compute the Counting function $F_{C}$ (with or without termination) with no knowledge about $\ell$, even when restricted to connected static simple networks with a known and arbitrarily small ratio $\ell/n$.
@@
Let us fix a positive integer $k$; we will construct an infinite class of networks whose ratio $\ell/n$ is $1/k$ as follows. For every $i\geq 3$, let $\mathcal{G}_{i}$ be the static network consisting of a cycle of $n_{i}=k\cdot i$ processes of which $\ell_{i}=i$ are leaders, such that the leaders are evenly spaced among the non-leaders. Assume that all processes get the same input (apart from their leader flags). Then, at any round, all the leaders in all of these networks have isomorphic views, which are independent of $i$. It follows that, if nothing is known about $\ell_{i}$ (other than the ratio $\ell_{i}/n_{i}$, which is fixed), all the leaders in all the networks $\mathcal{G}_{i}$ always give equal outputs. Since the number of processes $n_{i}$ depends on $i$, it follows that at most one of these networks can stabilize on the correct output $n_{i}$. ∎
Leaderless networks. We will now prove that the frequency-based functions introduced in Section 2 are the only functions that can be computed deterministically in anonymous networks without leaders. This result, together with Proposition 2.2, shows that the algorithms in Sections 4.2 and 4.3 can be generalized to all functions that can be computed in leaderless networks, i.e., the frequency-based functions.
Although the next theorem is implied by [30, Theorem III.1], here we provide an alternative and simpler proof.
@@
No function other than the frequency-based functions can be computed with no leader, even when restricted to connected static simple networks.
@@
Let $m_{1}$, $m_{2}$, …, $m_{k}$ be integers greater than $2$ with $\gcd(m_{1},m_{2},\dots,m_{k})=1$, and let $B$ be the complete $k$-partite graph with partite sets $V_{1}$, $V_{2}$, …, $V_{k}$ of sizes $m_{1}$, $m_{2}$, …, $m_{k}$, respectively. For any positive integer $\alpha$, construct the static network $G_{\alpha}$ consisting of $\alpha$ disjoint copies of $B$, augmented with $k$ cycles $C_{1}$, $C_{2}$, …, $C_{k}$ such that, for each $1\leq i\leq k$, the cycle $C_{i}$ spans all the $\alpha m_{i}$ processes in the $\alpha$ copies of $V_{i}$. Clearly, $G_{\alpha}$ is a connected static simple network.
Let the function $\lambda_{\alpha}$ assign input $z_{i}$ to all processes in the $\alpha$ copies of $V_{i}$ in $G_{\alpha}$, and let $\lambda=\lambda_{1}$. As a result, $\mu_{\lambda_{\alpha}}=\{(z_{1},\alpha m_{1}),(z_{2},\alpha m_{2}),\dots,(z_{k},\alpha m_{k})\}=\alpha\cdot\mu_{\lambda}$ for all $\alpha\geq 1$. Moreover, all the networks $G_{\alpha}$ have isomorphic history trees. This is because, at every round, each process in any of the copies of $V_{i}$ receives exactly two messages from other processes in copies of $V_{i}$ and exactly $m_{j}$ messages from processes in copies of $V_{j}$, for all $j\neq i$. Thus, it can be proved by induction that all processes in the copies of $V_{i}$ have isomorphic views, regardless of $\alpha$.
Due to the fundamental theorem of history trees Theorem 3.1, all the processes with input $z_{i}$ must give the same output $\psi(z_{i},\mu_{\lambda})=\psi(z_{i},\mu_{\lambda_{\alpha}})=\psi(z_{i},\alpha\cdot\mu_{\lambda})$, regardless of $\alpha$. Hence, by definition, only frequency-based functions can be computed in these networks. ∎
The following result justifies the assumptions made in Section 4.3 that processes have knowledge of an upper bound on $n$ or on the dynamic diameter $d$ of the network.
@@
No algorithm can solve the leaderless Average Consensus problem with explicit termination if nothing is known about the size of the network or its diameter, even when restricted to connected static simple networks.
@@
Assume for a contradiction that there is such an algorithm $\mathcal{A}$. Let $\mathcal{G}$ be a static network consisting of three processes forming a cycle, and assign input 00 to all of them. If the processes execute $\mathcal{A}$, they eventually output the mean value 00 and terminate, say in $t$ rounds.
Now construct a static network $\mathcal{G}^{\prime}$ consisting of a cycle of $22t+$ 2 processes $p_{1}$, $p_{2}$, …, $p_{2t+2}$; assign input $1$ to $p_{1}$ and input 00 to all other processes. It is easy to see that, from round 00 to round $t$, the view of the process $p_{t+1}$ is isomorphic to the view of any process in $\mathcal{G}$. Therefore, if $p_{t+1}$ executes $\mathcal{A}$, it terminates in $t$ rounds with the incorrect output 00. Thus, $\mathcal{A}$ is incorrect. ∎
@@ 6.2 Lower Bounds
Let $\mathcal{G}$ and $\mathcal{G}^{\prime}$ be two networks on $n$ and $n^{\prime}$ processes respectively, where $n\neq n^{\prime}$. Assume that there is a process $p$ in $\mathcal{G}$ and a process $p^{\prime}$ in $\mathcal{G}^{\prime}$ such that $p$ and $p^{\prime}$ have isomorphic views at round $t$. Then,
@@
Since $p$ and $p^{\prime}$ have isomorphic views at round $t$, they have isomorphic views at all rounds up to $t$. Thus, by Theorem 3.1, if $p$ and $p^{\prime}$ execute the same algorithm, they give equal outputs up to round $t$. Since the Counting function $F_{C}$ prescribes that $p$ must output $n$ and $p^{\prime}$ must output $n^{\prime}\neq n$, it is impossible for both processes to simultaneously give the correct output within $t$ rounds. In particular, no Counting algorithm can stabilize within $t$ rounds.
Moreover, if the execution of $p$ terminated within $t$ rounds, then the execution of $p^{\prime}$ would terminate at the same round, as well (again, due to Theorem 3.1). In that case, both processes would return the same output, which would be incorrect for at least one of them. In particular, no Counting algorithm can terminate within $t$ rounds in $\mathcal{G}$. ∎
@@
Let $(\mathcal{G}_{n})_{n\geq 1}$ be a sequence where $\mathcal{G}_{n}$ is a network on $n$ processes. Assume that there exists a constant $k>0$ such that, for infinitely many values of $n\in\mathbb{N}$, there is a process in $\mathcal{G}_{n}$ and a process in $\mathcal{G}_{n+k}$ that have isomorphic views at round $an+b$, for integer constants $a>0$ and $b$. Then, there is no algorithm that computes the Counting function $F_{C}$ in every $\mathcal{G}_{n}$ in less than $an-ak+b+1$ rounds. If termination is required, the bound improves to $an+b+1$ rounds.
@@
Let $m\in\mathbb{N}$ be such that $\mathcal{G}_{m}$ and $\mathcal{G}_{m+k}$ contain processes with isomorphic views at round $t=am+b$. Assume for a contradiction that a Counting algorithm stabilized in $f(n)\leq an-ak+b$ rounds in every $\mathcal{G}_{n}$. Then, it would stabilize in $f(m)\leq am-ak+b\leq t$ rounds in $\mathcal{G}_{m}$ and in $f(m+k)\leq a(m+k)-ak+b=t$ rounds in $\mathcal{G}_{m+k}$. This contradicts Lemma 6.5, which states that no Counting algorithm stabilizes within $t$ rounds in both networks.
@@
We first introduce a family of 1-union-connected dynamic networks. For any $n\geq 1$, we consider the dynamic network $\mathcal{G}_{n}$ whose topology at round $t$ is the graph $G^{(n)}_{t}$ defined on the system $\{p_{1},p_{2},\dots,p_{n}\}$ as follows. If $t\geq n-2$, then $G^{(n)}_{t}$ is the path graph $P_{n}$ spanning all processes $p_{1}$, $p_{2}$, …, $p_{n}$ in order. If $31\leq t\leq n-$ 3, then $G^{(n)}_{t}$ is $P_{n}$ with the addition of the single edge $\{p_{t+1},p_{n}\}$. We assume $p_{1}$ to be the leader and all other processes to have the same input.
@@
In 1-union-connected dynamic simple networks with a unique leader, no algorithm can compute the Counting function $F_{C}$ in less than $62n-$ 6 rounds. If termination is required, the bound improves to $42n-$ 4 rounds.
@@
Let us consider the network $\mathcal{G}_{n}$ as defined above. It is straightforward to prove by induction that, at every round $t\leq n-3$, the process $p_{t+1}$ gets disambiguated, while all processes $p_{t+2}$, $p_{t+3}$, …, $p_{n}$ are still indistinguishable. So, the history tree of $\mathcal{G}_{n}$ has a very regular structure, which is illustrated in Figures 9 and 10. By comparing the history trees of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$, we see that the leaders of the two systems have identical views up to round $52n-$ 5. Our claim now follows immediately from Corollary 6.6, with $k=1$, $a=2$, and $b=-5$
A slightly better lower bound can be obtained if we allow self-loops in the network. Consider the static network $\mathcal{G}_{n}$ consisting of a path graph spanning all $n$ processes, where one endpoint of the path is the unique leader and the other endpoint has a self-loop. It is easy to see that the leaders in $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ have identical views up to round $22n-$ 2. This implies a lower bound of $32n-$ 3 rounds for stabilization and $12n-$ 1 rounds for termination in static networks with self-loops and a unique leader.
Figure 9: The first rounds of the dynamic network $\mathcal{G}_{n}$ used in Proposition 6.7 (left) and the corresponding levels of its history tree (right), where $n=6$; the process in blue is the leader. The white nodes and the dashed edges in the history tree are not in the view of the leader at round $7$. The labels $p_{1}$, …, $p_{6}$ have been added for the reader’s convenience, and mark the processes that get disambiguated, as well as their corresponding nodes of the history tree, which have anonymity $1$.
Figure 10: The first rounds of the dynamic network $\mathcal{G}_{n+1}$ with $n=6$. Observe that the view of the leader at round $7$ is identical to the view highlighted in Figure 9. The intuitive reason is that, from round $1$ to round $n-3$, both networks have a cycle whose processes are all indistinguishable (and are therefore represented by a single node in the history tree), except for the one process with degree $3$. Thus, the history trees of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ are identical up to level $n-3$. After that, the two networks get disambiguated, but this information takes another $n-3$ rounds to reach the leader. Therefore, if the leader of $\mathcal{G}_{n}$ and the leader of $\mathcal{G}_{n+1}$ execute the same algorithm, they must have the same internal state up to round $52n-$ 5, due to Theorem 3.1. In particular, they cannot give different outputs up to that round, which leads to our lower bounds on stabilization and termination for the Counting problem.
@@
For any $\ell\geq 1$, no algorithm can compute the Counting function $F_{C}$ in all simple $\tau$-union-connected networks with $\ell$ leaders in less than $\tau(2n-\ell-5)$ rounds, which improves to $\tau(2n-\ell-3)$ rounds if termination is required.
@@
As shown in Proposition 6.7, there is a family of simple 1-union-connected networks $\mathcal{G}_{n}$, with $n\geq 1$, with the following properties. $\mathcal{G}_{n}$ has $\ell=1$ leader and $n$ processes in total; moreover, up to round $52n-$ 5, the leaders of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ have isomorphic views.
Let us fix $\ell\geq 1$, and let us construct $\mathcal{G}^{\prime}_{n}$, for $n\geq\ell$, by attaching a chain of $\ell-1$ additional leaders $p_{1}$, $p_{2}$, …, $p_{\ell-1}$ to the single leader $p_{\ell}$ of $\mathcal{G}_{n-\ell+1}$ at every round. Note that $\mathcal{G}^{\prime}_{n}$ has $n$ processes in total and a stable subpath $(p_{1},p_{2},\dots,p_{\ell})$ which is attached to the rest of the network via $p_{\ell}$.
It is straightforward to see that the process $p_{\ell}$ in $\mathcal{G}^{\prime}_{n}$ and the process $p_{\ell}$ in $\mathcal{G}^{\prime}_{n+1}$, which correspond to the leaders of $\mathcal{G}_{n-\ell+1}$ and $\mathcal{G}_{n-\ell+2}$ respectively, have isomorphic views up to round $32(n-\ell)-$ 3. Since the view of $p_{1}$ is completely determined by the view of $p_{\ell}$, and it takes $\ell-1$ rounds for any information to travel from $p_{\ell}$ to $p_{1}$, we conclude that the process $p_{1}$ in $\mathcal{G}^{\prime}_{n}$ and the process $p_{1}$ in $\mathcal{G}^{\prime}_{n+1}$ have isomorphic views up to round $42n-\ell-$ 4.
@@
Leaderless networks. Finally, we can use Proposition 6.7 to obtain a lower bound for the Average Consensus problem in leaderless networks. This lower bound shows that the leading term in the running time of the algorithm in Section 4.2, i.e., $2\tau n$, is optimal. The same holds for the algorithm in Section 4.3 assuming that $N=n$.
@@
No algorithm can solve the Average Consensus problem (with or without termination) in all simple $\tau$-union-connected leaderless networks in less than $\tau(2n-6)$ rounds, which improves to $\tau(2n-4)$ rounds if termination is required.
@@
According to Proposition 6.7, the number of processes $n$ in a network with $\ell=1$ and $\tau=1$ cannot be determined in less than $62n-$ 6 rounds ($42n-$ 4 rounds if termination is required). We can reduce this problem to Average Consensus with $\ell=0$ and $\tau=1$ as follows. In any given network with $\ell=\tau=1$, assign input $1$ to the leader and clear its leader flag; assign input 00 to all other processes. If the processes can compute the mean input value, $1/n$, they can invert it to obtain $n$ in the same number of rounds. It follows that Average Consensus with $\ell=0$ and $\tau=1$ cannot be solved in less than $62n-$ 6 rounds ($42n-$ 4 rounds if termination is required); this immediately generalizes to an arbitrary $\tau$ by Proposition 2.4. ∎

### v8

@@ 6 Negative Results
In this section we collect several negative results and counterexamples, which not only provide lower bounds asymptotically matching our algorithms’ running times, but also justify all of the assumptions made in Sections 4 and 5 about the a-priori knowledge of agents. Although some of these facts were previously known, here we offer simple and self-contained proofs based on history trees.
We point out that all the counterexamples provided in this section are simple networks; that is, all of our main results hold equivalently for networks modeled as multigraphs, as well as simple graphs. Moreover, all the counterexamples in Section 6.1 are static networks, implying that our impossibility results hold for dynamic, as well as static networks.
@@ 6.1 Unsolvable Problems
Showing that certain problems are not solvable in certain network models allows us to argue that our algorithms are universal, i.e., they can be applied to all solvable problems.
Networks with leaders. We will prove that the multiset-based functions introduced in Section 2 are the only functions that can be computed deterministically in anonymous networks with leaders, even when restricted to connected static simply networks. This result, together with Proposition 2.1, shows that the algorithms in Sections 4.4 and 5 can be generalized to all functions that can be computed in networks with leaders, i.e., the multiset-based functions.
@@
For any $\ell\geq 1$, no algorithm computes a non-multiset-based function (with or without termination) on all static simple networks in $\bigcup_{n\geq\ell}\mathcal{N}^{1}_{n,\ell}$.
@@
For $n\geq\ell$, let us consider the (static) network whose topology at round $t$ is the complete graph $G_{t}=K_{n}$, i.e., each agent receives messages from all other agents at every round. We can prove by induction that all nodes of the history tree other than the root have exactly one child. This is because any two agents with the same input always receive equal multisets of messages, and are therefore always indistinguishable. Thus, the history tree is completely determined by the multiset $\mu_{\lambda}$ of all agents’ inputs; moreover, an agent’s vista at any given round only depends on the agent’s own input and on $\mu_{\lambda}$ (the fact that $\ell$ is a constant does not affect this). By the fundamental theorem of history trees Theorem 3.1, this is enough to conclude that if an agent’s output stabilizes, that output must be a function of the agent’s own input and of $\mu_{\lambda}$, which is the defining condition of a multiset-based function. ∎
The following result justifies the assumption made in Sections 4.4 and 5 that agents have a-priori knowledge of the number of leaders $\ell$ in the system. It states that no algorithm can compute the Counting function $F_{C}$ (which, as we recall, is a multiset-based function) with no exact knowledge of $\ell$, even when restricted to connected static simple networks with a known and arbitrarily small ratio $\ell/n$.
@@
For any integer $k\geq 1$, no algorithm computes the Counting function $F_{C}$ (with or without termination) on all static simple networks in $\bigcup_{\ell\geq 1}\mathcal{N}^{1}_{k\ell,\ell}$.
@@
Let us fix $k\geq 1$; we will construct an infinite class of networks with $k\ell$ agents and $\ell$ leaders as follows. For every $i\geq 3$, let $\mathcal{G}_{i}$ be the static network consisting of a cycle of $n_{i}=k\cdot i$ agents of which $\ell_{i}=i$ are leaders, such that the leaders are evenly spaced among the non-leaders. Assume that all agents get the same input (apart from their leader flags). Then, at any round, all the leaders in all of these networks have isomorphic vistas, which are independent of $i$. It follows that, since nothing is known about $\ell_{i}$ (other than the ratio $\ell_{i}/n_{i}$, which is fixed), all the leaders in all the networks $\mathcal{G}_{i}$ always give equal outputs. Since the number of agents $n_{i}$ depends on $i$, at most one of these networks can stabilize on the correct output $n_{i}$. ∎
Leaderless networks. We will now prove that the frequency-based functions introduced in Section 2 are the only functions that can be computed deterministically in anonymous networks without leaders, even when restricted to connected static simple networks. This result, together with Proposition 2.2, shows that the algorithms in Sections 4.2 and 4.3 can be generalized to all functions that can be computed in leaderless networks, i.e., the frequency-based functions.
Although the next theorem is implied by [35, Theorem III.1], here we provide an alternative and simpler proof.
@@
No algorithm computes a non-frequency-based function (with or without termination) on all static simple networks in $\bigcup_{n\geq 1}\mathcal{N}^{1}_{n,0}$.
@@
Let $m_{1}$, $m_{2}$, …, $m_{k}$ be integers greater than $2$ with $\gcd(m_{1},m_{2},\dots,m_{k})=1$, and let $B$ be the complete $k$-partite graph with partite sets $V_{1}$, $V_{2}$, …, $V_{k}$ of sizes $m_{1}$, $m_{2}$, …, $m_{k}$, respectively. For any positive integer $\alpha$, construct the static network $G_{\alpha}$ consisting of $\alpha$ disjoint copies of $B$, augmented with $k$ cycles $C_{1}$, $C_{2}$, …, $C_{k}$ such that, for each $1\leq i\leq k$, the cycle $C_{i}$ spans all the $\alpha m_{i}$ agents in the $\alpha$ copies of $V_{i}$. Clearly, $G_{\alpha}$ is a connected static simple network.
Let the function $\lambda_{\alpha}$ assign input $z_{i}$ to all agents in the $\alpha$ copies of $V_{i}$ in $G_{\alpha}$, and let $\lambda=\lambda_{1}$. As a result, $\mu_{\lambda_{\alpha}}=\{(z_{1},\alpha m_{1}),(z_{2},\alpha m_{2}),\dots,(z_{k},\alpha m_{k})\}=\alpha\cdot\mu_{\lambda}$ for all $\alpha\geq 1$. Moreover, all the networks $G_{\alpha}$ have isomorphic history trees. This is because, at every round, each agent in any of the copies of $V_{i}$ receives exactly two messages from other agents in copies of $V_{i}$ and exactly $m_{j}$ messages from agents in copies of $V_{j}$, for all $j\neq i$. Thus, it can be proved by induction that all agents in the copies of $V_{i}$ have isomorphic vistas, regardless of $\alpha$.
Due to the fundamental theorem of history trees Theorem 3.1, all the agents with input $z_{i}$ must give the same output $\psi(z_{i},\mu_{\lambda})=\psi(z_{i},\mu_{\lambda_{\alpha}})=\psi(z_{i},\alpha\cdot\mu_{\lambda})$, regardless of $\alpha$. Hence, by definition, only frequency-based functions can be computed in these networks. ∎
The following result justifies the assumptions made in Section 4.3 that agents have knowledge of an upper bound on $n$ or on the dynamic diameter $d$ of the network. It states that, with no such knowledge, the leaderless Average Consensus problem (which, as we recall, is a frequency-based function) cannot be solved with explicit termination, even when restricted to connected static simple networks.
@@
No algorithm solves the Average Consensus problem with termination on all static simple networks in $\bigcup_{n\geq 1}\mathcal{N}^{1}_{n,0}$.
@@
Assume for a contradiction that there is such an algorithm $\mathcal{A}$. Let $\mathcal{G}$ be a static network consisting of three agents forming a cycle, and assign input 00 to all of them. If the agents execute $\mathcal{A}$, they eventually output the mean value 00 and terminate, say in $t$ rounds.
Now construct a static network $\mathcal{G}^{\prime}$ consisting of a cycle of $22t+$ 2 agents $p_{1}$, $p_{2}$, …, $p_{2t+2}$; assign input $1$ to $p_{1}$ and input 00 to all other agents. It is easy to see that, from round 00 to round $t$, the vista of the agent $p_{t+1}$ is isomorphic to the vista of any agent in $\mathcal{G}$. Therefore, if $p_{t+1}$ executes $\mathcal{A}$, it terminates in $t$ rounds with the incorrect output 00. Thus, $\mathcal{A}$ is incorrect. ∎
@@ 6.2 Lower Bounds
Let $\mathcal{G}$ and $\mathcal{G}^{\prime}$ be two networks on $n$ and $n^{\prime}$ agents respectively, where $n\neq n^{\prime}$. Assume that there is an agent $p$ in $\mathcal{G}$ and an agent $p^{\prime}$ in $\mathcal{G}^{\prime}$ such that $p$ and $p^{\prime}$ have isomorphic vistas at round $t$. Then,
@@
Since $p$ and $p^{\prime}$ have isomorphic vistas at round $t$, they have isomorphic vistas at all rounds up to $t$. Thus, by Theorem 3.1, if $p$ and $p^{\prime}$ execute the same algorithm, they give equal outputs up to round $t$. Since the Counting function $F_{C}$ prescribes that $p$ must output $n$ and $p^{\prime}$ must output $n^{\prime}\neq n$, it is impossible for both agents to simultaneously give the correct output within $t$ rounds. In particular, no Counting algorithm can stabilize within $t$ rounds.
Moreover, if the execution of $p$ terminated within $t$ rounds, then the execution of $p^{\prime}$ would terminate at the same round, as well (again, due to Theorem 3.1). In that case, both agents would return the same output, which would be incorrect for at least one of them. In particular, no Counting algorithm can terminate within $t$ rounds in $\mathcal{G}$. ∎
@@
Let $(\mathcal{G}_{n})_{n\geq 1}$ be a sequence where $\mathcal{G}_{n}$ is a network on $n$ agents. Assume that there exists a constant $k>0$ such that, for infinitely many values of $n\in\mathbb{N}$, there is an agent in $\mathcal{G}_{n}$ and an agent in $\mathcal{G}_{n+k}$ that have isomorphic vistas at round $an+b$, for integer constants $a>0$ and $b$. Then, there is no algorithm that computes the Counting function $F_{C}$ in every $\mathcal{G}_{n}$ in less than $an-ak+b+1$ rounds. If termination is required, the bound improves to $an+b+1$ rounds.
@@
Let $m\in\mathbb{N}$ be such that $\mathcal{G}_{m}$ and $\mathcal{G}_{m+k}$ contain agents with isomorphic vistas at round $t=am+b$. Assume for a contradiction that a Counting algorithm stabilized in $f(n)\leq an-ak+b$ rounds in every $\mathcal{G}_{n}$. Then, it would stabilize in $f(m)\leq am-ak+b\leq t$ rounds in $\mathcal{G}_{m}$ and in $f(m+k)\leq a(m+k)-ak+b=t$ rounds in $\mathcal{G}_{m+k}$. This contradicts Lemma 6.5, which states that no Counting algorithm stabilizes within $t$ rounds in both networks.
@@
We first introduce a family of 1-union-connected dynamic networks. For any $n\geq 1$, we consider the dynamic network $\mathcal{G}_{n}$ whose topology at round $t$ is the graph $G^{(n)}_{t}$ defined on the system $\{p_{1},p_{2},\dots,p_{n}\}$ as follows. If $t\geq n-2$, then $G^{(n)}_{t}$ is the path graph $P_{n}$ spanning all agents $p_{1}$, $p_{2}$, …, $p_{n}$ in order. If $31\leq t\leq n-$ 3, then $G^{(n)}_{t}$ is $P_{n}$ with the addition of the single edge $\{p_{t+1},p_{n}\}$. We assume $p_{1}$ to be the leader and all other agents to have the same input.
@@
No algorithm computes the Counting function $F_{C}$ on all simple networks in $\bigcup_{n\geq 1}\mathcal{N}^{1}_{n,1}$ in less than $62n-$ 6 rounds. If termination is required, the bound improves to $42n-$ 4 rounds.
@@
Let us consider the network $\mathcal{G}_{n}$ as defined above. It is straightforward to prove by induction that, at every round $t\leq n-3$, the agent $p_{t+1}$ gets disambiguated, while all agents $p_{t+2}$, $p_{t+3}$, …, $p_{n}$ are still indistinguishable. So, the history tree of $\mathcal{G}_{n}$ has a very regular structure, which is illustrated in Figures 9 and 10. By comparing the history trees of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$, we see that the leaders of the two systems have identical vistas up to round $52n-$ 5. Our claim now follows immediately from Corollary 6.6, with $k=1$, $a=2$, and $b=-5$
A slightly better lower bound can be obtained if we allow self-loops in the network. Consider the static network $\mathcal{G}_{n}$ consisting of a path graph spanning all $n$ agents, where one endpoint of the path is the unique leader and the other endpoint has a self-loop. It is easy to see that the leaders in $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ have identical vistas up to round $22n-$ 2. This implies a lower bound of $32n-$ 3 rounds for stabilization and $12n-$ 1 rounds for termination in static networks with self-loops and a unique leader.
Figure 9: The first rounds of the dynamic network $\mathcal{G}_{n}$ used in Proposition 6.7 (left) and the corresponding levels of its history tree (right), where $n=6$; the agent in blue is the leader. The white nodes and the dashed edges in the history tree are not in the vista of the leader at round $7$. The labels $p_{1}$, …, $p_{6}$ have been added for the reader’s convenience, and mark the agents that get disambiguated, as well as their corresponding nodes of the history tree, which have anonymity $1$.
Figure 10: The first rounds of the dynamic network $\mathcal{G}_{n+1}$ with $n=6$. Observe that the vista of the leader at round $7$ is identical to the vista highlighted in Figure 9. The intuitive reason is that, from round $1$ to round $n-3$, both networks have a cycle whose agents are all indistinguishable (and are therefore represented by a single node in the history tree), except for the one agent with degree $3$. Thus, the history trees of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ are identical up to level $n-3$. After that, the two networks get disambiguated, but this information takes another $n-3$ rounds to reach the leader. Therefore, if the leader of $\mathcal{G}_{n}$ and the leader of $\mathcal{G}_{n+1}$ execute the same algorithm, they must have the same internal state up to round $52n-$ 5, due to Theorem 3.1. In particular, they cannot give different outputs up to that round, which leads to our lower bounds on stabilization and termination for the Counting problem.
@@
For any $\tau\geq 1$ and $\ell\geq 1$, no algorithm computes the Counting function $F_{C}$ on all simple networks in $\bigcup_{n\geq\ell}\mathcal{N}^{\tau}_{n,\ell}$ in less than $\tau(2n-\ell-5)$ rounds. If termination is required, the bound improves to $\tau(2n-\ell-3)$ rounds.
@@
As shown in Proposition 6.7, there is a family of simple 1-union-connected networks $\mathcal{G}_{n}$, with $n\geq 1$, with the following properties. $\mathcal{G}_{n}$ has $\ell=1$ leader and $n$ agents in total; moreover, up to round $52n-$ 5, the leaders of $\mathcal{G}_{n}$ and $\mathcal{G}_{n+1}$ have isomorphic vistas.
Let us fix $\ell\geq 1$, and let us construct $\mathcal{G}^{\prime}_{n}$, for $n\geq\ell$, by attaching a chain of $\ell-1$ additional leaders $p_{1}$, $p_{2}$, …, $p_{\ell-1}$ to the single leader $p_{\ell}$ of $\mathcal{G}_{n-\ell+1}$ at every round. Note that $\mathcal{G}^{\prime}_{n}$ has $n$ agents in total and a stable subpath $(p_{1},p_{2},\dots,p_{\ell})$ which is attached to the rest of the network via $p_{\ell}$.
It is straightforward to see that the agent $p_{\ell}$ in $\mathcal{G}^{\prime}_{n}$ and the agent $p_{\ell}$ in $\mathcal{G}^{\prime}_{n+1}$, which correspond to the leaders of $\mathcal{G}_{n-\ell+1}$ and $\mathcal{G}_{n-\ell+2}$ respectively, have isomorphic vistas up to round $32(n-\ell)-$ 3. Since the vista of $p_{1}$ is completely determined by the vista of $p_{\ell}$, and it takes $\ell-1$ rounds for any information to travel from $p_{\ell}$ to $p_{1}$, we conclude that the agent $p_{1}$ in $\mathcal{G}^{\prime}_{n}$ and the agent $p_{1}$ in $\mathcal{G}^{\prime}_{n+1}$ have isomorphic vistas up to round $42n-\ell-$ 4.
@@
Leaderless networks. Finally, we can use Proposition 6.7 to obtain a lower bound for the Average Consensus problem in leaderless networks. This lower bound shows that the leading term in the running time of the algorithm in Section 4.2, i.e., $2\tau n$, is optimal. The same holds for the algorithm in Section 4.3, assuming that $N=n$.
@@
For any $\tau\geq 1$, no algorithm solves the Average Consensus problem on all simple networks in $\bigcup_{n\geq 1}\mathcal{N}^{\tau}_{n,0}$ in less than $\tau(2n-6)$ rounds. If termination is required, the bound improves to $\tau(2n-4)$ rounds.
@@
According to Proposition 6.7, the number of agents $n$ in a network with $\ell=1$ and $\tau=1$ cannot be determined in less than $62n-$ 6 rounds ($42n-$ 4 rounds if termination is required). We can reduce this problem to Average Consensus with $\ell=0$ and $\tau=1$ as follows. In any given network with $\ell=\tau=1$, assign input $1$ to the leader and clear its leader flag; assign input 00 to all other agents. If the agents can compute the mean input value, $1/n$, they can invert it to obtain $n$ in the same number of rounds. It follows that Average Consensus with $\ell=0$ and $\tau=1$ cannot be solved in less than $62n-$ 6 rounds ($42n-$ 4 rounds if termination is required); this immediately generalizes to an arbitrary $\tau$ by Proposition 2.5. ∎